Wrtie the point-slope form of the line that passes through (4,-1) with slope of 2.
my ans was: y + 1 =2(x - 4)is this close, or correct?

Answer by venugopalramana(3286)

Wrtie the point-slope form of the line that passes through (4,-1) with slope of 2.
my ans was: y + 1 =2(x - 4)is this close, or correct?
EXCELLENT KEEP IT UP 100% CORRET.THE FORMULA IS
Y-Y1=M(X-X1)...M=SLOPE=2.....X1=4...X COORDINATE OF POINT AND Y1=-1...Y COORDINATE OF POINT